# Revisiting Koehler’s experiment of measuring the ratio of the specific heats of air by self-sustained oscillations: a more concise theoretical interpretation

### v1

@@ Abstract
We revisit Koehler’s experiment, a clever modification of Rüchardt’s experiment designed to measure the ratio of specific heats of gas. The theory of self-sustained oscillations in Koehler’s experiment was provided by Koehler (1950). However, its lengthy and dense analysis may pose challenges to readers due to the complexity of the calculations. Following Koehler’s approximation for pressure changes, we explicitly present the model equations as piecewise linear differential systems and qualitatively analyze the periodic solutions from a geometric perspective. This concise and transparent approach addresses a fundamental question about Koehler’s experiment: why is the oscillation frequency nearly equal to the Rüchardt frequency? Our analysis avoids intricate calculations and will be particularly helpful for teachers and students who encounter Koehler’s experiment in general physics laboratory classes.
@@ I Introduction
Koehler’s experiment for measuring the ratio of specific heats $\gamma=C_{p}/C_{v}$ of air (or other gases) is a modification of Rüchardt’s experiment, a classic experiment widely used in general physics laboratory courses Koehler1951 ; Lange2000 . As shown in Fig. 1(a), the apparatus of Rüchardt’s experiment consists of a large vessel with a vertical tube containing a closely fitting steel ball. To facilitate the discussion that follows, it is useful to briefly recall the theory of Rüchardt’s experiment. When the ball is stationary at its equilibrium position, the equilibrium pressure satisfies the equation $P_{0}=B+\frac{mg}{\pi R^{2}}$, Where $B$ is atmospheric pressure, mm and $R$ are the mass and radius of the ball. When the ball is displaced from its equilibrium position and released, its rapid motion induces pressure fluctuations due to the adiabatic compression or expansion of the gas within the container. For the ball, we have
$m\ddot{x}=-\pi R^{2}\cdot\Delta P$ (1)
Where $x$ is the displacement of the ball from its equilibrium position, and $\Delta P$ is the gas pressure change in the vessel. Considering the radiabatic equation $PV^{\gamma}=const.$,
we have
$\Delta P=-\frac{\gamma P\Delta V}{V}\sim-\frac{\gamma P_{0}\Delta V}{V_{0}}=-\frac{\gamma P_{0}\cdot\pi R^{2}}{V_{0}},$ (2)
where $V_{0}$ is the volume of the vessel. Substituting it into Eq.(1), it can be seen that the motion of the ball is simple harmonic oscillation. The relationship between the period of oscillation and $\gamma$ is given by
@@
Although Rüchardt’s method has the advantages of simplicity and mechanical intuitiveness, it is not as simple as it first appears. Unavoidable frictional damping and gas leakage limit the number of oscillations, typically to only a few, preventing precise measurement of the period. In 1951, Koehler described an interesting modification involving a gas-feeding pump and introducing a small hole in the tube (as shown in Fig. 1(b))Koehler1951 . The ball is lifted by gas overpressure and falls back due to the pressure drop from gas leakage through the hole. By carefully adjusting the gas flow, the ball can oscillate symmetrically about the hole indefinitely. The oscillations closely approximate harmonic motion, and the relationship between the period and $\gamma$ is still described by Eq.(3). The advantages of Koehler’s method are evident: it allows prolonged measurement of the period, thereby improving the ease and precision of measurements. Therefore, compared to Rüchardt’s experiment, Koehler’s experiment offers greater advantages for teaching purposes.
Figure 1: (a) and (b) Schematic diagrams of Rüchardt apparatus and Koehler apparatus. (c) An image of experimental apparatus.
In China, many universities incorporate the Koehler experiment into general physics laboratory courses. However, there is limited evidence of its widespread adoption globally. A search of online resources reveals few instances of its use in English-speaking countries, particularly in American colleges and universities. Even in experimental demonstrations of the specific heat ratio of gases on YouTube, nearly all are based on Rüchardt’s design. By incorporating a simple air pump (costing less than \$10), the Koehler experiment achieves sustained oscillation through a clever design, enabling more precise measurement of the gas-specific heat ratio. Thus, we argue that the Koehler experiment provides significant pedagogical advantages over the Rüchardt experiment in student laboratory teaching and has the potential for broader adoption in other regions in the future. Furthermore, even in schools that currently use the Rüchardt experiment, instructors could easily adapt the setup to implement the Koehler version or allow students to explore it as an independent project, given its simplicity. For reference, the detailed parameters of the experimental apparatus employed in this study are provided in Appendix A.
@@
In courses using Koehler’s experiment, the theoretical introduction typically follows the framework of the original Rüchardt experiment. Students are generally informed that the ball’s motion approximates harmonic behavior, as described by Eq.(3). However, an issue arises: the supply and leakage of gas indicate that the pressure fluctuations within the container are not solely determined by the adiabatic changes in gas volume. The pressure change corresponding to the restoring force is no longer linearly proportional to the displacement of the ball. Why do these motions still closely approximate harmonic oscillations and satisfy Eq.(3)? In fact, the ball’s motion represents a nonlinear, self-sustained oscillation. Koehler presented a theoretical explanation for such self-sustained oscillations in another paper Koehler1950 . Assuming the ball’s motion to be sinusoidal, he used nonlinear dynamic equations to demonstrate that the pressure fluctuations are also approximately sinusoidal, with the oscillation frequency closely matching the Rüchardt frequency. However, this analysis is indirect, with a lengthy and dense derivation, potentially posing challenges to readers given the complexity of the calculations. To our knowledge, no other studies provide a more accessible theoretical analysis of this issue. This might explain why textbooks often mention it briefly, stating, for example, that “the pressure changes caused by gas supply and leakage are relatively small, so…” without offering detailed explanations.
@@
The following sections of this paper will revisit this issue and provide a concise theoretical interpretation of the self-sustained oscillations. The paper is organized as follows. In section 2, following Koehler’s approximation for pressure changes, we explicitly present the dynamic equations. In section 3, we dimensionless the model equations and expressed them as 3D piecewise linear differential systems, and analyzed the behavior of the systems near their fixed points in a 2D projection plane. In section 4, we qualitatively analyze the properties of periodic solutions of the system from a geometric perspective, demonstrating that the oscillation frequency in Koehler’s experiment is nearly equal to the Rüchardt frequency.
II Dynamic equations of self-sustained oscillations
@@
Let $x$ and $v$ denote the vertical coordinate and velocity of the ball, $P$ the pressure in the vessel (assumed spatially uniform) which is a function of the ball’s coordinate $x$ and time $t$. The dynamic equations are
@@
Eq.(4c) shows that the rate of pressure change consists of two parts: the term ${\frac{\partial P}{\partial x}}{\frac{dx}{dt}}$ caused by the adiabatic change of gas volume due to the motion of the ball, and the term $\frac{\partial P}{\partial t}$ contributed by gas supply and leakage. For the adiabatic process of an ideal gas,
@@
For $\frac{\partial P}{\partial t}$, here we use the approximation from Koehler’s paperKoehler1950 . The contribution to $\frac{\partial P}{\partial t}$ due to gas supply is approximated as a positive constant
$\frac{\partial P}{\partial t}_{supply}=jB.$ (6)
The contribution to $\frac{\partial P}{\partial t}$ due to the gas leakage is approximated as proportional to the difference in pressure between the inside and outside of the container.
$\frac{\partial P}{\partial t}_{leakage}=-g(x)(P-B),\quad g(x)=\begin{cases}e+f,&\quad x\geq 0\\ e,&\quad x<0\end{cases},$ (7)
The coordinate origin is defined at the center of the small hole on the tube. When $x<0$, gas escapes solely through the annular gap between the ball and the tube wall. When $x\geq 0$, gas escapes through both the annular gap around the ball and the small hole. The finite size of the small hole is neglected, resulting in a discontinuity of $g(x)$ at 0. However, this effect is minor and not crucial to the main point. By substituting Eqs. (5), (6), and (7) into Eqs. (4), the final dynamic equations are derived as
@@
$\frac{dP}{dt}=-\frac{\gamma P_{0}\pi R^{2}}{V_{0}}v+jB-g(x)(P-B).$ (8c)
When the term jBg(x)(PB)jB-g(x)(P-B) is removed, the equations reduce to those describing harmonic motion in the ideal Rüchardt experiment. The Rüchardt frequency is given by $\omega^{2}=\frac{\gamma P_{0}\pi^{2}R^{4}}{mV_{0}}$(equivalent to Eq. (3)). The term involving the piecewise function $g(x)$ introduces nonlinearity. As a result, the equations can no longer be solved analytically.
@@ III Preliminary qualitative analysis
where $\tau=\omega t,y=\frac{x}{R},u=\frac{v}{R\omega},p=\frac{P-P_{0}}{B},p_{0}=\frac{P_{0}}{B},A=\frac{\pi RB}{m\omega^{2}},J=\frac{j}{\omega}$, and
@@
In mathematics, Eqs.(9) can be classified as piecewise linear differential systems (PWLD systems). These systems consist of two linear subsystems separated by the plane defined by $\{y=0\}$, referred to as the separation plane. They can be written in matrix form as
$\dot{\mathbf{x}}=\begin{cases}M_{+}\mathbf{x}+\mathbf{e}_{+}&\mathrm{if}\quad y\geq 0\\ M_{-}\mathbf{x}+\mathbf{e}_{-}&\mathrm{if}\quad y<0\end{cases}$ (10)
@@
$\mathbf{x}=\begin{pmatrix}y\\ u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{-}\end{pmatrix}$
@@
PWLD systems are widely used to model real-world phenomena in fields such as electrical circuits, control systems, impact and friction mechanics, and biological systems. However, analyzing periodic solutions, which often correspond to steady-state responses in general PWLD systems, is notoriously challenging Jeffrey2018 . This difficulty arises because PWLD systems autonomously switch their behavior at specific boundaries, leading to discontinuities and distinct sets of governing equations in different regions of the phase space. Accurately handling transitions across these boundaries often requires solving implicit or transcendental equations, which can be particularly challenging or even infeasible in practice. Despite these challenges, the following sections of this paper will provide an approximate analysis to demonstrate if the periodic solutions exist, specifically, solutions in which the small ball passes through the hole position once upward and once downward during a single period, then the system’s motion frequency is indeed very close to the Rüchardt frequency.
Analyzing fixed points (equilibrium points) yields valuable insights for dynamical systems. In both half-space $\{y\geq 0\}$ and $\{y<0\}$, Eqs.(9) do not exhibit global fixed points. However, note that on the right-hand side of the equations, the variable $y$ appears only within $G(y)$, whose constant value is piecewise determined by $y$. Thus, $y$ can be treated as a parameter, allowing for an analysis of the system in the $(u,p)$ projection plane for each half-space, $\{y\geq 0\}$ and $\{y<0\}$.
@@
where $G$ represents $G(y)$, with its values in the two half-space $\{y\geq 0\}$ and $\{y<0\}$ denoted as $G_{+}$ and $G_{-}$, respectively. Let $\frac{du}{d\tau}=0,\frac{dp}{d\tau}=0$, we obtain the fixed point of this subsystem, denoted as $(u^{*},p^{*})$, given by
$u^{*}=A\left[J-G(p_{0}-1)\right],\quad p^{*}=0.$
Next, the local behavior around the fixed point is analyzed. Let $u=u^{*}+\eta(\tau),p=p^{*}+\xi(\tau)$, where $\eta(\tau)$ and $\xi(\tau)$ represent small deviations from the fixed point. Substituting these into Eqs. (11), the system is rewritten as
$\frac{d}{d\tau}\binom{\eta}{\xi}=\begin{bmatrix}0&A\\ -\frac{1}{A}&-G\end{bmatrix}\binom{\eta}{\xi}$. (12)
The Jacobian matrix $\begin{bmatrix}0&A\\ -\frac{1}{A}&-G\end{bmatrix}$ has eigenvalues $\lambda=-\frac{G}{2}\pm i\sqrt{1-\frac{G^{2}}{4}}$. Since $G$ is positive and much smaller than $1$ (as shown in Appendix B, which estimates the experimental values of parameters such as $A$, $J$, and $G$ ), the term $1-\frac{G^{2}}{4}$ is positive. Therefore, the fixed point $(u^{*},p^{*})$ is a stable focus, and any trajectory in the plane will spiral towards this point. The perturbations $\eta(\tau)$ and $\xi(\tau)$ can be expressed as
@@
where $C_{1}$, $\Phi$ are undetermined coefficients related to initial values, and $tan(\varphi)=\frac{\sqrt{1-\frac{G^{2}}{4}}}{\frac{G}{2}}$. These results show that the perturbations in the ball’s velocity and the gas pressure are both characterized by decaying oscillations with a frequency of $\sqrt{1-\frac{G^{2}}{4}}\sim 1$. This indicates that in Koehler’s experiment, whether the ball is below the small hole (in half space y<0{y<0}) or above it (in half space $y\geq 0$), the changes in the ball’s velocity and the gas pressure remain closely approximated by harmonic motion, with a frequency nearly identical to the Rüchardt frequency $\omega$.
It is important to note that, so far, we can only state that the oscillation frequency of the system within each subspace is close to the Rüchardt frequency. However, we cannot arbitrarily say that the frequency of the overall motion, obtained by combining the trajectories from each subspace, is also close to the Rüchardt frequency. This is because the exact duration for which the system remains in each subspace is not known. For instance, let $T_{0}$ represent the Rüchardt period. If the system spends $\frac{T_{0}}{2}$ in one half-space and $\frac{T_{0}}{3}$ in another half-space, which may be possible, the total period of the motion will be $(\frac{T_{0}}{2}+\frac{T_{0}}{3})$, not $T_{0}$, resulting in a frequency that deviates from the Rüchardt frequency. In the following section, we will present a geometric analysis to demonstrate that, under appropriate experimental conditions, when the system alternates between two half-spaces, the overall motion period approaches $T_{0}$, and the oscillation frequency remains close to the Rüchardt frequency.
IV Discussion on Periodic Solutions of Model Equations
In this section, we will estimate the time the system spends in each subspace from a geometric perspective in phase space and demonstrate that the overall motion period approaches $T_{0}$. We find that the original 3D PWLD system with one separation plane $\{y=0\}$ can be equivalently simplified to a 2D PWLD system with a single separation straight line, which is more analytically tractable. Although this section is somewhat lengthy, it essentially describes how the system’s trajectory revolves around two alternating focal points in a periodic manner.

### v2

@@ Abstract
We revisit Koehler’s experiment, a clever modification of Rüchardt’s experiment designed to measure the ratio of specific heats of gas. However, the lengthy and dense analysis shared by Koehler in his 1950 paper may pose challenges to readers due to the complexity of the calculations. Following Koehler’s approximation for pressure changes, we explicitly present the model equations as piecewise linear differential systems and qualitatively analyze the periodic solutions from a geometric perspective. This concise and transparent approach addresses a fundamental question about Koehler’s experiment: why is the oscillation frequency nearly equal to the Rüchardt frequency? Our analysis avoids intricate calculations and should help educators introduce Koehler’s experiment in general physics laboratory classes.
@@ I Introduction
Koehler’s experiment for measuring the ratio of specific heats $\gamma=C_{p}/C_{v}$ of air (or other gases) is a modification of Rüchardt’s experiment, a classic experiment widely used in general physics laboratory courses Koehler1951 ; Lange2000 . As shown in Fig. 1(a), the apparatus for Rüchardt’s experiment consists of a large sealed vessel connected to a vertical tube, in which a closely fitting steel ball can move freely. When the ball is stationary at its equilibrium position, the pressure in the vessel satisfies the equation
$P_{0}=P_{\mathrm{atm}}+\frac{mg}{\pi R^{2}},$ (1)
where $P_{\mathrm{atm}}$ is the atmospheric pressure, and mm and $R$ are the mass and radius of the ball, respectively. When the ball is displaced vertically from equilibrium and released, it undergoes rapid oscillatory motion along the $x$ axis (see Fig. 1(a)), inducing pressure fluctuations through the adiabatic compression and expansion of the gas in the vessel. The equation of motion for the ball can be written as
$m\ddot{x}=\pi R^{2}\,\Delta P.$ (2)
In this equation, $x$ is the displacement of the ball from its equilibrium position (positive upward), and $\Delta P$ is the pressure change in the vessel. Assuming an adiabatic process, the gas satisfies
$PV^{\gamma}=\mathrm{const.}~.$
Because the volume change induced by the motion of the ball is small compared with the total volume, the adiabatic relation can be linearized about equilibrium, yielding
$\Delta P=-\frac{\gamma P_{0}\Delta V}{V_{0}}=-\frac{\gamma P_{0}\cdot\pi R^{2}x}{V_{0}},$ (3)
where $V_{0}$ is the total gas volume when the ball is at equilibrium, including the vessel and a portion of the tube. In practice, $V_{0}$ can be approximated by the volume of the vessel, since the contribution from the tube is negligible. In this limit, the ball executes simple harmonic motion. The relationship between the period of oscillation $T$ and $\gamma$ is given by
@@
Although Rüchardt’s method has the advantages of simplicity and mechanical intuitiveness, it is not as simple as it first appears. Due to frictional damping and gas leakage, the ball executes only a few oscillations before coming to rest, preventing precise measurement of the period. In 1951, Koehler described an interesting modification of this method, involving a gas-feeding pump and a small hole in the tube (as shown in Fig. 1(b))Koehler1951 . The ball is lifted by gas overpressure and, once its vertical position is above that of the hole, falls back as gas leaks. When the gas flow is adjusted to be within a certain range, the ball can oscillate symmetrically about the hole indefinitely. The oscillations closely approximate harmonic motion, with a period still given by Eq.(4). The advantages of Koehler’s method are evident: it allows prolonged measurement operation, thereby improving the ease and precision of the measurements. Therefore, compared to Rüchardt’s experiment, Koehler’s experiment offers greater advantages for teaching purposes, although it appears to be less widely spread. To encourage educators to explore this method in their course, the detailed parameters of the experimental apparatus employed in this study are provided in Appendix A.
Figure 1: (a) and (b) Schematic diagrams of Rüchardt apparatus and Koehler apparatus.
@@
In courses using Koehler’s experiment, the theoretical introduction typically follows the framework of the original Rüchardt experiment. Students are generally informed that the ball’s motion approximates harmonic behavior, as described by Eq.(4). However, the derivation for Eq.(4) cannot be identically replicated, since the pressure fluctuations are not only the result of the adiabatic changes in the gas volume, but are also due to the incoming gas and to the gas leakage through the hole. Koehler presented a theoretical explanation for the ball’s self-sustained oscillations in another paper Koehler1950 . Assuming the ball’s motion to be sinusoidal, the author used nonlinear dynamical equations to demonstrate that the pressure fluctuations are also approximately sinusoidal, with a frequency closely matching the Rüchardt frequency. However, this analysis is indirect and lengthy, and might be challenging to readers. To our knowledge, no other studies provide a more accessible theoretical analysis of this issue.
@@
The following sections of this paper will revisit this issue and provide a concise theoretical interpretation of the self-sustained oscillations. The paper is organized as follows. In section II, following Koehler’s approximation for pressure changes, we present the dynamical equations governing the ball’s motion. In section III, we nondimensionalize the model equations and express them as a three-dimensional piecewise linear differential system. The behavior near the fixed points is then analyzed in a two-dimensional projection. In section  IV, we qualitatively analyze the properties of the system’s periodic solutions from a geometric perspective, demonstrating that the oscillation frequency in Koehler’s experiment is nearly equal to the Rüchardt frequency.
II Dynamical equations of self-sustained oscillations
@@
Let $x$ and $v$ denote the vertical coordinate and velocity of the ball, respectively. Let $P$ denote the pressure in the vessel, which is a function of the ball’s coordinate $x$ and time $t$, and is assumed to be uniform in the vessel. Let $P_{0}$ denote the equilibrium pressure, given by Eq. (1). The dynamical equations are
@@
Eq.(5c) shows that the rate of pressure change consists of two parts: the fist term arising from the adiabatic change of the gas volume due to the ball’s motion, and the term $\partial P/\partial t$ due to the gas supply and leakage. For the adiabatic process of an ideal gas,
@@
Here also, we have assumed that the ball’s displacement induces a negligible change in the vessel’s volume: $P/V\sim P_{0}/V_{0}$. For $\partial P/\partial t$, we use the approximation from Koehler’s paperKoehler1950 .
The contribution to $\partial P/\partial t$ due to the gas supply (an aquarium air pump) is approximated as a positive constant
$\left(\frac{\partial P}{\partial t}\right)_{\mathrm{supply}}=jP_{\mathrm{atm}}.$ (7)
This approximation is reasonable because the air pump used here operates approximately as a constant-flow source over the relevant pressure range. Assuming a constant volumetric flow rate $Q$, the ideal gas equation
$PV=nRT$
gives
$\frac{dP}{dt}=\frac{RT}{V}\frac{dn}{dt},$
Since the molar inflow rate satisfies
$\frac{dn}{dt}=\frac{P_{\mathrm{atm}}Q}{RT},$
one obtains
$\frac{dP}{dt}=\frac{P_{\mathrm{atm}}Q}{V}.$
Therefore,
$j=\frac{Q}{V}.$
For the pump used in the experiment described in the Appendices, the maximum flow rate is approximately
$Q\sim 7~\mathrm{L/min}\sim 1.2\times 10^{-4}~\mathrm{m^{3}\,s^{-1}}.$
For a typical container volume $V=2.5~\mathrm{L}$, the corresponding order of magnitude of $j$ is estimated to be
$j\sim 0.05~\mathrm{s^{-1}}.$
Moreover, the characteristic timescale associated with the pressure increase, $1/j\sim 20~\mathrm{s}$, is typically much longer than the oscillation period of the system, which is of order one second. Hence, over several oscillation cycles, the pressure increase produced by the pump may be regarded as approximately linear in time, which justifies the approximation used in Eq. (7).
Similarly, since the pressure variation induced by the gas leakage is also small, the contribution to $\partial P/\partial t$ due to leakage is approximated as being proportional to the pressure difference between inside and outside the container:
$\left(\frac{\partial P}{\partial t}\right)_{\mathrm{leakage}}=-g(x)(P-P_{\mathrm{atm}}),\quad g(x)=\begin{cases}e+f,&\quad x\geq 0\\ e,&\quad x<0\end{cases},$ (8)
where the origin of the $x$-axis is defined at the center of the small hole on the tube. When $x<0$, the gas escapes solely through the unavoidable annular gap between the ball and the tube wall. When $x\geq 0$, the gas escapes through both the annular gap around the ball and the small hole. The finite size of the small hole is neglected, resulting in a discontinuity of $g(x)$ at 0. However, this effect is minor and not crucial to the main point. By substituting Eqs. (6), (7), and (8) into Eqs. (5), the final dynamical equations are derived as
@@
$\frac{dP}{dt}=-\frac{\gamma P_{0}\pi R^{2}}{V_{0}}v+jP_{\mathrm{atm}}-g(x)(P-P_{\mathrm{atm}}).$ (9c)
When the term $jP_{\mathrm{atm}}-g(x)(P-P_{\mathrm{atm}})$ is removed, the equations reduce to those describing the harmonic motion in the ideal Rüchardt experiment. The Rüchardt frequency is given by $\omega^{2}=\pi^{2}\gamma P_{0}R^{4}/(mV_{0})$(equivalent to Eq. (4)). The term involving the piecewise function $g(x)$ introduces nonlinearity. As a result, the equations can no longer be solved analytically.
@@ III Preliminary qualitative analysis
where $\tau=\omega t,y=x/R,u=v/(R\omega),p=(P-P_{0})/P_{\mathrm{atm}},p_{0}=P_{0}/P_{\mathrm{atm}},A=(\pi RP_{\mathrm{atm}})/(m\omega^{2}),J=j/\omega$, and
@@
Equation (10) is a piecewise linear differential system (PWLD systems). This system consists of two linear subsystems separated by the plane defined by $\{y=0\}$, referred to as the separation plane. They can be written in matrix form as
$\dot{\mathbf{z}}=\begin{cases}M_{+}\mathbf{z}+\mathbf{e}_{+}&\mathrm{if}\quad y\geq 0\\ M_{-}\mathbf{z}+\mathbf{e}_{-}&\mathrm{if}\quad y<0\end{cases}$ (11)
@@
$\mathbf{z}=\begin{pmatrix}y\\ u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{-}\end{pmatrix}$
@@
PWLD systemsLlibre2018 are widely used to model real-world phenomena in fields such as electrical circuits, control systems, impact and friction mechanics, and biological systems. However, analyzing periodic solutions, which often correspond to steady-state responses in general PWLD systems, is notoriously challenging Jeffrey2018 . This difficulty arises because PWLD systems switch their behavior at the separation plane, leading to discontinuities and distinct sets of governing equations in different regions of the phase space. Accurately handling transitions across these boundaries often requires solving implicit or transcendental equations, which can be particularly challenging or even infeasible in practice. Despite these challenges, the following sections provide an approximate analysis showing that, if periodic solutions exist—i.e., solutions for which the ball passes through $x=0$ at regular time intervals—the oscillation frequency is very close to the Rüchardt frequency.
Analyzing fixed points (equilibrium points) yields valuable insights for dynamical systems. In both half-spaces $\{y\geq 0\}$ and $\{y<0\}$, Eqs.(10) do not exhibit global fixed points. However, note that on the right-hand side of the equations, the variable $y$ appears only within $G(y)$, whose constant value is piecewise determined by $y$. Thus, $y$ can be treated as a parameter, allowing for an analysis of the system in the $(u,p)$ projection plane for each half-space, $\{y\geq 0\}$ and $\{y<0\}$:
@@
where $G\equiv G(y)$. Setting $du/d\tau=0$ and $dp/d\tau=0$, we obtain the fixed point of this subsystem, denoted by $(u^{*},p^{*})$, as
$u^{*}=A\left[J-G(p_{0}-1)\right],\quad p^{*}=0.$ (13)
To analyze the local behavior near the fixed point, we introduce small perturbations
$u=u^{*}+\eta(\tau),\quad p=p^{*}+\xi(\tau),$
where $\eta(\tau)$ and $\xi(\tau)$ are small deviations. Substituting into Eqs. (12), we obtain
$\frac{d}{d\tau}\begin{pmatrix}\eta\\ \xi\end{pmatrix}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G\end{pmatrix}\begin{pmatrix}\eta\\ \xi\end{pmatrix}$. (14)
The Jacobian matrix in Eq. (14) has eigenvalues
$\lambda=-\frac{G}{2}\pm i\sqrt{1-\frac{G^{2}}{4}}.$
Since $G$ is positive and much smaller than $1$ (as shown in Appendix B, which estimates the experimental values of parameters such as $A$, $J$, and $G$ ), the term $1-G^{2}/4$ is positive. Therefore, the fixed point $(u^{*},p^{*})$ is a stable focus, and any trajectory in the plane will spiral towards this point. The perturbations $\eta(\tau)$ and $\xi(\tau)$ can be expressed as
@@
where $C_{1}$, $\Phi$ are undetermined coefficients related to initial values, and
$tan(\varphi)=\frac{\sqrt{1-\frac{G^{2}}{4}}}{\frac{G}{2}}.$
These results show that the perturbations in the ball’s velocity and the gas pressure are both characterized by decaying oscillations with a frequency of $\sqrt{1-G^{2}/4}\sim 1$. This indicates that in Koehler’s experiment, whether the ball is below the small hole (in the half space y<0{y<0}) or above it (in the half space $y\geq 0$), the changes in the ball’s velocity and in the gas pressure remain closely approximated by harmonic motion, with a frequency nearly identical to the Rüchardt frequency $\omega$.
It is important to note that, so far, we can only state that the oscillation frequency of the system within each subspace is close to the Rüchardt frequency. However, we cannot arbitrarily say that the frequency of the overall motion, obtained by combining the trajectories from each subspace, is also close to the Rüchardt frequency. This is because the exact duration for which the system remains in each subspace is not known. For instance, let $T_{0}$ represent the Rüchardt period. If the system spends $T_{0}/2$ in one half-space and $T_{0}/3$ in another half-space, which may be possible, the total period of the motion will be $(T_{0}/2+T_{0}/3)$, not $T_{0}$, resulting in a frequency that deviates from the Rüchardt frequency. In the following section, we present a geometric analysis to demonstrate that, under appropriate experimental conditions, when the system alternates between the two half-spaces, the overall motion period approaches $T_{0}$, and the oscillation frequency remains close to the Rüchardt frequency.
IV Periodic Solutions of Model Equations
In this section, we estimate the time the system spends in each subspace using geometric considerations in the phase space and show that the overall period of motion approaches $T_{0}$. We show that the original three-dimensional (3D) PWLD system with a single separation plane $\{y=0\}$ can be equivalently simplified to a two-dimensional (2D) PWLD system with a single separation line, which is more analytically tractable. Although this section is somewhat lengthy, it essentially describes how the system’s trajectory revolves around two alternating focal points in a periodic manner.